OpenStudy (lukecrayonz):

Let’s suppose you throw a ball straight up with an initial speed of 50 feet per second from a height of 6 feet. Find the parametric equations that describe the motion of the ball as a function of time. How long is the ball in the air? Determine when the ball is at maximum height. Find its maximum height.

OpenStudy (lukecrayonz):

Satellite, you were on this really fast.. :P

OpenStudy (anonymous):

usual formula is \[g(t)=-16t^2+v_0t+h_0\] here \[v_0=50,h_0=6\]

OpenStudy (anonymous):

i meant \(h(t)\)

OpenStudy (anonymous):

in any case \[h(t)=-16t^2+50t+6\] for your example. to see how long it is in the air, set \(h(t)=0\) and solve for \(t\)

OpenStudy (lukecrayonz):

I just saw how my textbook wanted me to do it, i'll show you how it wanted.

OpenStudy (anonymous):

max height is at the vertex of this parabola, which opens down if your text wants to you take a limit, tell it to go pound sand. no point in it vertex of the parabola is where \(t=-\frac{b}{2a}\)

OpenStudy (lukecrayonz):

x=v[0]cos(theta))t+h y=-(1/2)gt^2+(v[0]sintheta)t+h

OpenStudy (lukecrayonz):

http://screensnapr.com/v/FpJu8X.png

OpenStudy (lukecrayonz):

Well they just use different equations from you, whats the difference though?

OpenStudy (anonymous):

ball in this case is thrown straight up, so \(\theta =90, \cos(90)=0,\sin(90)=1\)

OpenStudy (lukecrayonz):

I like your way better :P